On 15 October 2017, Desmond bought 5 000 shares in Red Ltd at a cost of R5 per share. On 25 October 2017, Red Ltd was taken over by Blue plc. The shareholders of Red Ltd received 3 shares in Blue plc for every 2 shares in Red Ltd. On 25 October 2017 shares in Blue plc were valued at R3.50 each. How many shares does Desmond’s hold in Blue plc at 25 October 2017 and how much did they cost?

7 500 shares at a cost of R26 250

7 500 shares at a cost of cost R25 000

5 000 shares at a cost R25 000

5000 shares at a cost of R17 500
Answer #1

7 500 shares at a cost of R26 250

Number of shares:
15th October 5000 shares
25th October 5000/2*3=7500 shares

Total Cost:
15th October =5000*5=25000
25th October =7500*3.5=26250.00